The Express BanquetExpress Banquet has two restaurants that are open 24 hours a day. Fixed costs...

The

Express BanquetExpress Banquet

has two restaurants that are open 24 hours a day. Fixed costs for the two restaurants together total

$450,000 per year. Service varies from a cup of coffee to full meals. The average sales check per customer is

$ 8.00.The average cost of food and other variable costs for each customer is $ 3.20 The income tax rate is

30​%. Target net income is $105,000.

Requirements

1. Compute the revenues needed to earn the target net income.

2. How many customers are needed to break​ even? To earn net income of $105,000​?

3. Compute net income if the number of customers is 165,000.
Answer #1

Answer:

1) Statement showing computations
Particulars Amount($)
average sales check per customer                            8.00
average cost of food and other variable costs                            3.20
Contribution Margin                            4.80
CM Ratio = 4.80/8.00 60%
Target net income                       1,05,000
Income before tax = 105,000/(1-.30)                     1,50,000
Fixed costs                     4,50,000
Desired contribution = 150000+450000                     6,00,000
Desired Sales = 600,000/60%                   10,00,000
2) No of customer to break even = 450000/4.80                        93,750
customers needed to earn net income of $105000 = 600,000/4.80                     1,25,000
3) If customers are 165,000 = 165000*4.80 - 450000                     3,42,000